### v2

# BranchGRPO: Stable and Efficient GRPO with Structured Branching in Diffusion Models

Abstract
Recent progress in aligning image and video generative models with Group Relative Policy Optimization (GRPO) has improved human preference alignment, yet existing approaches still suffer from high computational cost due to sequential rollouts and large numbers of SDE sampling steps, as well as training instability caused by sparse rewards. In this paper, we present BranchGRPO, a method that restructures the rollout process into a branching tree, where shared prefixes amortize computation and pruning removes low-value paths and redundant depths. BranchGRPO introduces three contributions: (1) a branch sampling scheme that reduces rollout cost by reusing common segments; (2) a tree-based advantage estimator that converts sparse terminal rewards into dense, step-level signals; and (3) pruning strategies that accelerate convergence while preserving exploration. On HPDv2.1 image alignment, BranchGRPO improves alignment scores by up to 16% over strong baselines, while reducing per-iteration training time by nearly 55%. On WanX-1.3B video generation, it further achieves higher Video-Align scores with sharper and temporally consistent frames compared to DanceGRPO.
1 Introduction
Diffusion and flow-matching models have advanced image and video generation with high fidelity, diversity, and controllability (Ho et al., 2020; Lipman et al., 2022; Liu et al., 2022). However, large-scale pretraining alone cannot ensure alignment with human intent, as outputs often miss aesthetic, semantic, or temporal expectations. Reinforcement learning from human feedback (RLHF) addresses this gap by directly adapting models toward human-preferred outcomes (Ouyang et al., 2022).
Within RLHF, Group Relative Policy Optimization (GRPO) has shown strong stability and scalability across text-to-image and text-to-video tasks (Liu et al., 2025a; Xue et al., 2025). However, when applied to diffusion and flow-matching models, current GRPO variants still face two fundamental bottlenecks: (1) Inefficiency. Standard GRPO adopts a sequential rollout design, where each trajectory must be independently sampled under both the old and new policies. This incurs $O(N\cdot T)$ complexity with denoising steps $T$ and group size $N$, leading to significant computational redundancy and limiting scalability in large-scale image and video generation tasks. (2) Sparse rewards. Most methods compute only a single reward from the final denoised image or video and propagate it uniformly across all denoising steps. This trajectory-level supervision neglects informative signals contained in intermediate states, resulting in unstable optimization, inaccurate credit assignment, and high-variance gradients. This raises a central challenge: how can we accurately attribute sparse outcome rewards to the specific denoising steps that truly contribute to high-quality generations?
To overcome these limitations, we propose BranchGRPO, a tree-structured policy optimization framework for diffusion and flow models. BranchGRPO replaces inefficient independent sequential rollouts with a branching structure, where scheduled split steps in the denoising process allow each trajectory to stochastically expand into multiple sub-trajectories while reusing shared prefixes. This design amortizes computation across common segments and aligns naturally with the stepwise nature of denoising, substantially improving sampling efficiency while reducing computational cost.
Crucially, the tree structure enables a novel reward fusion and depth-wise advantage estimation mechanism. Instead of assigning a single terminal reward to every denoising step, BranchGRPO aggregates rewards from leaf nodes predicted by the reward model and propagates them backward through depth-wise normalization. This produces layer-wise advantages that allow for finer-grained credit assignment and more stable policy optimization. In addition, BranchGRPO integrates width- and depth-level pruning strategies that remove low-value or redundant branches. These pruning mechanisms accelerate training, reduce computational overhead, and enhance effective exploration by reallocating resources to promising regions of the trajectory space.
Our contributions are threefold:
  • We introduce BranchGRPO, a tree-structured rollout framework that replaces independent sequential sampling with branching trajectories. By inserting split points during denoising, trajectories expand into multiple sub-trajectories while sharing early prefixes, which reduces redundancy and improves exploration efficiency.
  • We propose a new reward fusion and depth-wise advantage estimation method that converts sparse terminal rewards into dense, step-level signals, yielding more stable optimization.
  • We design complementary width- and depth-pruning strategies that lower backpropagation cost and further improve alignment.
  • We validate the effectiveness of BranchGRPO on both text-to-image and image-to-video alignment tasks, demonstrating its general applicability across modalities.
[…]
2 Related Work
Diffusion models (Ho et al., 2020; Rombach et al., 2022) and flow matching models (Lipman et al., 2022; Liu et al., 2022) have become dominant paradigms for visual generation due to their strong theoretical foundations and ability to generate high-quality content efficiently. While pretraining establishes the generative prior, aligning outputs with nuanced human preferences requires reinforcement learning from human feedback (RLHF). In natural language processing, RLHF has proven highly successful for aligning large language models (LLMs) (Ouyang et al., 2022; Christiano et al., 2017), where methods such as PPO and GRPO enable stable preference-driven post-training. These successes have inspired adaptation of RLHF to vision.
In the visual domain, RLHF for diffusion has been developed along two main directions. Reward-model-based approaches such as ImageReward (Xu et al., 2023) backpropagate learned rewards through the denoising process. Direct Preference Optimization (DPO) (Rafailov et al., 2023) has also been extended to diffusion, leading to Diffusion-DPO (Wallace et al., 2024) and Videodpo (Liu et al., 2025c), which achieve competitive alignment without explicit reward modeling. Policy-gradient formulations such as DDPO (black2023ddpo) and DPOK (Fan et al., 2023) further explore online optimization but often face stability challenges. Meanwhile, standardized reward models including HPS-v2.1 (Wu et al., 2023), VideoAlign (Liu et al., 2025b) enable systematic comparison of alignment algorithms on image and video tasks.
More recently, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has been introduced as a scalable alternative to PPO for preference optimization. DanceGRPO (Xue et al., 2025) pioneers the application of GRPO to visual generation, unifying diffusion and flow models via SDE reformulation and demonstrating stable optimization across text-to-image, text-to-video, and image-to-video tasks. Flow-GRPO (Liu et al., 2025a) extends GRPO to flow matching models by introducing ODE-to-SDE conversion and denoising reduction for improved efficiency. MixGRPO (Li et al., 2025a) further enhances efficiency with a mixed ODE–SDE sliding-window scheme, though it still faces trade-offs between overhead and performance. Our work continues this line by introducing BranchGRPO, which leverages branching rollouts, depth-wise reward fusion, and structured pruning to improve both stability and efficiency; while related to TreePO in LLMs (Li et al., 2025b), our method adapts tree-structured rollouts specifically to diffusion dynamics.
3 Method
3.1 Does Branch Rollout Harm Diversity?
[…]
3.2 Branch Rollout Algorithm
Preliminaries.
Given a prompt, BranchGRPO reformulates denoising into a tree-structured process. We align terminology with tree search: (i) depth $T$ denotes the number of denoising steps; (ii) width $w$ is the number of completed trajectories (leaves); (iii) branching steps $\mathcal{B}$ indicate split timesteps; (iv) branch correlation $s$ controls the correlation among child noises; (v) branching factor $K$ is the number of children per split.
Branch sampling.
Unlike prior GRPO variants such as DanceGRPO and FlowGRPO that rely on sequential rollouts, where each trajectory is sampled independently from start to finish, BranchGRPO reorganizes the process into a tree-structured rollout (Figure 1). For each prompt, we initialize a root node with a same initial noise $z_{0}\sim\mathcal{N}(0,I)$ and then denoise step by step along the reverse SDE. At designated split steps $\mathcal{B}$, the current state expands into $K$ children, producing multiple sub-trajectories that share early prefixes but diverge afterward. The branching is achieved by injecting stochastic perturbations into the SDE transition, with a hyperparameter $s$ controlling the correlation strength among child noises. This mechanism balances exploration diversity and stability while keeping the marginal distribution unchanged. The rollout continues until reaching the maximum depth $T$, at which point all leaves are collected for reward evaluation. Formally, the baseline reverse SDE is
$\mathrm{d}z_{t}=\Big{(}f(z_{t},t)-\tfrac{1}{2}g^{2}(t)\nabla\log p_{t}(z_{t})\Big{)}\mathrm{d}t\;+\;g(t)\,\mathrm{d}w_{t},$ (1)
while our branching modification at split step $i\in\mathcal{B}$ produces
$z_{i+1}^{(b)}=\mu_{\theta}(z_{i},t_{i})+s_{i}\sqrt{h_{i}}\,\xi_{b},\qquad\xi_{b}=\frac{\xi_{0}+s\,\delta\xi_{b}}{\sqrt{1+s^{2}}},\quad b=1,\dots,K,$ (2)
where $\xi_{0},\delta\xi_{b}\sim\mathcal{N}(0,I)$ and $s\!\geq\!0$ adjusts inter-branch correlation. When $s=0$, all branches collapse to a single path, while as $s\!\to\!\infty$, the branches become independent.Intuitively, smaller $s$ yields highly correlated children (stable but limited exploration), whereas larger $s$ produces more independent trajectories (diverse but noisier).
Figure 3: Left: branching rollouts generate diverse leaves. Middle: leaf rewards are fused upward by path-probability weighting. Right: depth-wise normalization and pruning yield dense advantages and reduce cost.
3.3 Reward Fusion and Depth-wise Advantage Estimation
The branch rollout constructs a full trajectory tree in which all leaf nodes share common ancestral prefixes. This shared structure makes it natural to represent the rewards of intermediate nodes through their descendant leaves, analogous to how genetic traits of ancestors can be inferred from their offspring.
In contrast, existing GRPO variants update every denoising step using only the final reward, which ignores informative intermediate states and leads to unstable credit assignment with high variance (see Fig. 3). BranchGRPO leverages its tree structure to overcome this limitation: leaf rewards are propagated upward to internal nodes, normalized layer by layer, and transformed into dense step-wise advantages—without introducing any additional models or computational overhead. Our optimization objective remains a clipped GRPO loss defined on tree edges:
$J(\theta)=\mathbb{E}\Bigg{[}\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\min\big{(}\rho_{e}(\theta)A(e),\ \mathrm{clip}(\rho_{e}(\theta),1-\epsilon,1+\epsilon)A(e)\big{)}\Bigg{]},$ (3)
where $\mathcal{E}$ denotes tree edges and $\rho_{e}(\theta)$ is the importance ratio.
Reward fusion.
Directly averaging leaf rewards to estimate internal node values can be dominated by rare but noisy trajectories, leading to unstable gradients. To mitigate this, we fuse descendant rewards using path-probability weighting: let $r_{\ell}$ be the reward of a leaf $\ell$. For an internal node $n$ with children $\mathcal{L}(n)$, we fuse its rewards by
$\bar{r}(n)=\sum_{\ell\in\mathcal{L}(n)}w_{\ell}^{(n)}\,r_{\ell},\qquad w_{\ell}^{(n)}=\frac{\exp(\beta\,s_{\ell})}{\sum_{j\in\mathcal{L}(n)}\exp(\beta\,s_{j})},$ (4)
where $s_{\ell}$ is the path log-probability under the behavior policy and $\beta$ controls the strength of weighting. When $\beta=0$, this reduces to simple averaging; larger $\beta$ emphasizes high-likelihood paths and reduces the influence of noisy, low-probability leaves, in all our experiments we set $\beta=1.0$.
We consider two alternatives for aggregating leaf rewards. Uniform averaging treats all leaves equally is robust to log-prob calibration errors and encourages exploration by preserving contributions from low-probability leaves. However, it deviates from the expected return under the behavior policy and can introduce high variance when the number of branches is large.
In contrast, path-weighted scheme aligns node-level credit assignment with the behavior distribution, typically reducing variance and yielding more stable updates. Yet, it relies on log-prob calibration and may overly concentrate weight on high-likelihood paths in deep trees. We empirically compare both variants (Sec. 4.4), showing that path weighting improves stability in practice, while uniform averaging provides stronger exploratory diversity.
Depth-wise normalization.
The tree depth corresponds to denoising steps, and nodes at the same depth share the same noise level. This makes rewards across nodes within a depth directly comparable, while rewards across different depths may differ drastically in scale due to varying noise states. To ensure fair comparison, we normalize aggregated rewards separately within each depth $d$:
$A_{d}(n)=\frac{\bar{r}(n)-\mu_{d}}{\sigma_{d}+\epsilon},\qquad\mu_{d}=\mathrm{mean}_{n\in\mathcal{N}_{d}}\bar{r}(n),\ \ \sigma_{d}=\mathrm{std}_{n\in\mathcal{N}_{d}}\bar{r}(n),$ (5)
where $\mathcal{N}_{d}$ denotes the set of nodes at depth $d$. Each edge advantage $A(e)$ is then inherited from its child node and optionally clipped to $A_{\max},A_{\max}]$. This per-step standardization aligns gradient magnitudes with the denoising process and prevents deeper nodes from being dominated by shallow ones.
Compared to GRPO’s prompt-level normalization, which uniformly broadcasts a single terminal reward, our branch-based scheme yields process-dense credit signals that respect the denoising trajectory. By normalizing rewards at the step level, BranchGRPO produces more stable gradients and more accurate credit assignment to the intermediate denoising steps that truly matter.
3.4 Pruning Strategies
While branch rollouts improve efficiency and provide dense process rewards, an excessive number of branches may induce exponential growth in trajectory count, leading to prohibitive backpropagation cost. Recent work (Li et al., 2025a) has revealed that standard GRPO training contains substantial redundancy and pruning potential. Motivated by this, we introduce two complementary pruning strategies in the context of BranchGRPO: width pruning, which reduces the number of leaves used for gradient updates, and depth pruning, which skips unnecessary denoising layers.
Importantly, pruning is applied only after reward fusion and depth-wise normalization, and affects backpropagation but not forward rollouts or reward evaluation. This design ensures that all trajectories contribute to reward signals, while gradients are computed only for the selected subset of nodes.
Width Pruning.
After computing rewards and advantages for all leaves $\mathcal{L}$, we optionally select a subset of them for gradient updates. We study two simple modes. The first is Parent-Top1, which retains the higher-reward child of each parent in the last branching step. This strategy halves gradient cost while ensuring coverage of all parents, yielding stable but slightly less diverse updates. The second is Extreme selection, which preserves both the globally best and worst $b$ leaves by reward score. By explicitly maintaining strong positive and negative signals, this mode injects more exploratory feedback but introduces higher variance.
Depth Pruning.
Notably, recent work MixGRPO (Li et al., 2025a) achieves strong performance by hybrid ODE–SDE, effectively acting as a form of depth pruning. This motivates our explicit exploration of depth pruning within BranchGRPO.
Specifically, we maintain a set of pruned depths $\mathcal{D}$ and ignore gradients from nodes at these layers while still performing full forward rollouts for reward evaluation. To avoid permanently excluding certain steps, we employ a sliding window: the pruned depths shift gradually toward deeper layers as training progresses, until reaching a predefined maximum depth. Formally, pruning is enabled in all of training steps, and every fixed interval the pruning window slides one step deeper, until reaching a pre-defined stop depth.
This dynamic scheme improves efficiency while still accounting for the contributions of gradients from different depths, thus balancing computational savings with policy effectiveness.
Algorithm 1 BranchGRPO Training Process
1:dataset $\mathcal{C}$; policy $\pi_{\theta}$; behavior policy $\pi_{\theta_{\text{old}}}$; reward models $\{R_{k}\}$; denoising steps $T$; branching steps $\mathcal{B}$; branching factor $K$; branch correlation $s$
2:for iteration $m=1$ to $M$ do
3:  $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
4:  Sample batch $\mathcal{C}_{b}\subset\mathcal{C}$
5:  for prompt $c\in\mathcal{C}_{b}$ do
6:   Sample root noise $z_{0}\sim\mathcal{N}(0,I)$
7:   Build rollout tree $\mathcal{T}$ with $\pi_{\theta_{\text{old}}}$:
8:   for $t=T$ to 0 do
9:     if $t\in\mathcal{B}$ then
10:      Branch into $K$ children with correlation $s$
11:     else
12:      Single-step denoising
13:     end if
14:   end for
15:   Evaluate rewards for leaves $\mathcal{L}(\mathcal{T})$
16:   Reward fusion: aggregate leaf rewards upward (path-prob. weights)
17:   Depth-wise normalization: standardize per depth, assign edge advantages
18:   Pruning (optional): select nodes for backprop only
19:   Compute clipped GRPO loss $J(\theta)$
20:  end for
21:  Update policy: gradient ascent on $J(\theta)$
22:end for
4 Experiments
4.1 Experiment Setup
We evaluate BranchGRPO on HPDv2.1 (Wu et al., 2023) (103k training and 400 balanced test prompts). The backbone is FLUX.1-Dev, with rollouts augmented by branching and process-dense rewards; we also study a pruned variant (BranchGRPO-Pruning). Baselines include DanceGRPO and MixGRPO under identical settings. We report efficiency (NFE, iteration time, GPU-hours) and quality (HPS-v2.1, PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023)).
4.2 Implementation Details
We set the tree depth to $d=20$ and the branch factor to $b=2$, yielding $16$ leaves per rollout before pruning. The branching steps $\mathcal{B}$ use three presets: Dense $(0,3,6,9)$ as the default, Mixed $(0,5,15,20)$, and Sparse $(0,10,13,19)$. The branch correlation sweeps $s\in\{0,1,2,4,8\}$. Training runs for 300 optimizer steps with gradient accumulation $g=12$ and per-GPU batch size =2=2, on $16\times$ NVIDIA H20z GPUs. Optimization uses AdamW (learning rate $1\times 10^{-5}$, weight decay $1\times 10^{-4}$) with bf16 precision and EMA weights stored on CPU. All GRPO-related hyperparameters are kept identical across methods, with full details deferred to the supplementary material.
Figure 4: Qualitative comparison of generations from Flux, DanceGRPO, and our BranchGRPO.
[…]
FLUX - - - 0.313 0.227 1.112
DanceGRPO(tf=1.0) 20 20 698 0.360 0.229 1.189
DanceGRPO(tf=0.6) 20 12 469 0.353 0.228 1.219
MixGRPO (20,5) 20 5 289 0.359 0.228 1.211
BranchGRPO 13.68 13.68 493 0.363 0.229 1.233
BranchGRPO-WidPru 13.68 8.625 314 0.364 0.230 1.300
BranchGRPO-DepPru 13.68 8.625 314 0.369 0.231 1.319
Table 1: Efficiency–quality comparison under identical data, rewards, and optimization schedules. Best and second-best per column are in bold and underline. Speedup is relative to DanceGRPO (tf=1.0). NFE counts the expected number of denoising steps where the policy is updated; for branching methods this is an effective NFE that reflects split density and shared prefixes.
4.3 Main Results
Table 1 compares efficiency and alignment quality across methods, while Figure 1 shows reward dynamics during training. BranchGRPO achieves consistently better alignment than DanceGRPO while using substantially fewer function evaluations and shorter iteration times. MixGRPO improves efficiency but at the cost of weaker alignment, whereas BranchGRPO provides a more favorable balance, and pruning variants further reduce computation without hurting performance.
The reward curves in Figure 1 also highlight the stability of our approach. DanceGRPO with reduced timestep fractions suffers from strong reward fluctuations, while the full-timestep variant converges more smoothly but at much higher cost. BranchGRPO combines the best of both, offering faster reward growth and more stable trajectories. Figure 4 further illustrates qualitative improvements: compared with Flux and DanceGRPO, our method produces more faithful, detailed, and coherent generations across diverse prompts.
Taken together, these results demonstrate that BranchGRPO is both more efficient and more stable than existing GRPO variants, yielding superior preference alignment under significantly reduced computational overhead.
4.4 Ablation Studies
We conduct a series of ablation studies to better understand the design choices in BranchGRPO. Unless otherwise stated, all experiments are carried out under the same training setup as in Section 4.3. The following analyses highlight how different branching configurations and aggregation strategies affect efficiency, reward quality, and stability.
Branch Correlation.
Figure 5(a) shows the effect of varying the branch correlation $s$. Smaller values ($s=1.0,2.0$) limit exploration and lead to slower reward growth, while very large values ($s=8.0$) destabilize early training. A moderate setting ($s=4.0$) achieves the best trade-off, reaching the highest reward and stable convergence, confirming that stochastic branching is necessary but should be carefully tuned.
Branching Steps.
We next vary the positions of split timesteps (Figure 5(b)). Early splits such as $(0,3,6,9)$ promote faster reward increase in the early stage, whereas later splits like $(9,12,15,18)$ delay exploration and yield lower rewards. Intermediate schedules (e.g., $(3,6,9,12)$) balance efficiency and reward quality, suggesting that early splits are generally more effective for exploration.
Branch Density.
Finally, we compare different densities of split points while keeping the overall horizon fixed (Figure 5(c)). Although all configurations eventually converge to similar reward levels, denser splits (e.g., $(0,3,6,9)$) accelerate early training, while sparser configurations converge more slowly. This indicates that increasing the density of branching in the early phase improves sample efficiency without harming stability.
Figure 5: Ablation on branching hyperparameters without pruning. (a) Moderate correlation (s=4s{=}4) best balances exploration and stability. (b) Early split schedules yield higher rewards than late splits. (c) Denser early branching accelerates reward growth while converging to similar final levels.
Figure 6: Ablations on pruning and reward fusion. Left: pruning policies (none, depth, width-Parent-Top1, width-Extreme-$b$). Right: reward-fusion strategies (uniform vs. path-weighted, with/without width pruning). Depth pruning achieves the best final reward, while path-weighted fusion improves stability across settings.
Path-weighted vs. Uniform Fusion.
Figure 6 (right) compares uniform averaging with path-probability weighting under identical training settings. Uniform averaging shows higher variance and a clear late-stage plateau, whereas path-weighted fusion delivers consistently higher and more stable rewards throughout training. This empirically supports Sec. 3.3: uniform averaging is exploration-friendly but noisy, while path weighting aligns credit assignment with the behavior distribution and improves convergence at no extra cost.
Pruning.
Figure 6 (left) evaluates pruning applied after depth-wise normalization and only for backprop. Depth pruning (sliding window over denoising steps) achieves the best final reward, consistently surpassing the no-pruning baseline by reallocating compute to informative early steps while preserving forward sampling diversity. Width pruning (Parent-Top1) halves gradient updates and yields the smoothest, lowest-variance curve, but its final reward is slightly below depth pruning. Width pruning (Extreme-$b$) keeps the globally best and worst $b$ leaves; it injects strong positive/negative signals and remains competitive at the end, but exhibits higher variance than the other settings.
4.5 Video Generation Results
We further evaluate BranchGRPO on video generation using WanX-1.3B I2V (Wan et al., 2025). We compare three settings: the base model, DanceGRPO, and BranchGRPO. Figure 7 presents qualitative frames and quantitative reward curves.
Qualitatively, the base model exhibits severe temporal flickering, while DanceGRPO partially improves consistency but still struggles with detail preservation. BranchGRPO produces sharper frames with coherent objects across time.
Quantitatively, BranchGRPO achieves faster convergence and higher rewards, measured by the Video-Align motion quality metric (Liu et al., 2025b), confirming the benefit of branching rollouts for video generation.
Figure 7: Video generation results on WanX-1.3B I2V. Left: qualitative frame comparisons across three settings. (a) Without GRPO: severe temporal flickering and deformation. (b) DanceGRPO: improved but still inconsistent and blurry details. (c) BranchGRPO: sharper and temporally coherent frames. Right: reward curves showing faster convergence and higher final rewards with BranchGRPO compared to DanceGRPO.
5 Conclusion
We introduced BranchGRPO, a reinforcement learning framework that mitigates reward sparsity and rollout inefficiency in diffusion and flow models. By reformulating rollouts as tree-structured trajectories with reward fusion and lightweight pruning, BranchGRPO achieves denser credit assignment, faster convergence, and higher alignment quality at lower cost. These results establish BranchGRPO as a stable, efficient alternative to existing GRPO variants and a step toward scalable preference optimization for complex generative tasks.
Appendix A Appendix
This appendix provides additional theoretical proofs, implementation details, and experimental results that complement the main paper. The contents are organized as follows:
  • Section B: Hyperparameter settings used in all experiments.
  • Section C: Theoretical analysis, including proofs for branch noise construction, reward fusion, and variance reduction.
  • Section D: Additional experiments, including ablations, more text-to-image and image-to-video results.
  • Failure Cases: Qualitative examples where BranchGRPO fails, highlighting current limitations.
Appendix B Hyperparameter Settings
Table 2 summarizes the detailed hyperparameter configuration used in our experiments. All hyperparameters are kept identical across all methods, including DanceGRPO and MixGRPO, to ensure a fair comparison.
Table 2: Hyperparameter settings used in all experiments.
Parameter Value Parameter Value
Random seed 42 Learning rate $1\times 10^{-5}$
Train batch size 2 Weight decay $1\times 10^{-4}$
SP size 1 Mixed precision bfloat16
SP batch size 2 Grad. checkpointing Enabled
Dataloader workers 4 Max grad norm 0.01
Grad. accum. steps 12 Warmup steps 0
Checkpoint steps 40 Use TF32 Yes
Resolution $720\times 720$ Sampling steps 16
Eta 0.3 Sampler seed 1223627
Num. generations 12 Shift (branch offset) 3
Use group reward Yes Ignore last step Yes
Clip range $1\times 10^{-4}$ Adv. clip max 5.0
Use EMA Yes EMA decay 0.995
Init same noise Yes
Appendix C Theoretical Analysis (Detailed Proofs)
C.1 Branch Noise Construction and Boundary Distribution Preservation
Consider the reverse SDE discretized by Euler–Maruyama:
$z_{i+1}=\mu_{\theta}(z_{i},t_{i})+g_{i}\,\eta_{i},\qquad\eta_{i}\sim\mathcal{N}(0,I),\ \ g_{i}:=g(t_{i})\sqrt{h_{i}}.$ (6)
At a split step $i\in B$, we construct $K$ branch noises as
$\xi_{b}=\frac{\xi_{0}+s\,\delta\xi_{b}}{\sqrt{1+s^{2}}},\qquad\xi_{0},\delta\xi_{b}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,I),$ (7)
so that $\xi_{b}\sim\mathcal{N}(0,I)$ for each $b$, and $\mathrm{Cov}(\xi_{b},\xi_{b^{\prime}})=\tfrac{1}{1+s^{2}}I$ for $b\neq b^{\prime}$.
Each child branch then updates as
$z^{(b)}_{i+1}=\mu_{\theta}(z_{i},t_{i})+g_{i}\,\xi_{b}.$ (8)
Lemma 1 (Single-step marginal preservation).
For any fixed parent $z_{i}$, we have
$z^{(b)}_{i+1}\ \stackrel{{\scriptstyle d}}{{=}}\ \mu_{\theta}(z_{i},t_{i})+g_{i}\eta_{i},\qquad\eta_{i}\sim\mathcal{N}(0,I).$
Lemma 2 (Leaf marginal preservation).
Conditioned on prefix $(z_{0},\eta_{0},\dots,\eta_{i-1})$, each branch $z^{(b)}_{N}$ generated by the split rule has the same distribution as a baseline SDE sample $z_{N}$.
Theorem 1 (Boundary distribution invariance).
Under the branching construction above, for any set of split steps $B$, each leaf $z^{(b)}_{N}$ has the same marginal law as a baseline SDE rollout. Hence, branching does not alter the final generator distribution.
C.2 Reward Fusion: Unbiasedness and Variance Reduction
Let $L(n)$ be the leaf set of a node $n$. Each leaf has reward $r_{\ell}=r(z_{N}^{(\ell)})$. Define the conditional expected return
$V(n):=\mathbb{E}[\,r(z_{N})\mid n\,].$
Uniform fusion.
$\bar{r}(n)=\frac{1}{|L(n)|}\sum_{\ell\in L(n)}r_{\ell}.$ (9)
Then
$\mathbb{E}[\bar{r}(n)\mid n]=V(n),\qquad\mathrm{Var}(\bar{r}(n)\mid n)=\frac{\sigma^{2}(n)}{|L(n)|},$
where $\sigma^{2}(n)=\mathrm{Var}(r_{\ell}\mid n)$.
Path-probability weighted fusion.
If leaves are drawn from proposal $q(\ell\mid n)$ with weights
$w_{\ell}=\frac{p_{\text{beh}}(\ell\mid n)}{q(\ell\mid n)},$
then the IS estimator
$\hat{r}_{\mathrm{IS}}(n)=\frac{1}{|L(n)|}\sum_{\ell\in L(n)}w_{\ell}r_{\ell}$
is unbiased: $\mathbb{E}[\hat{r}_{\mathrm{IS}}(n)\mid n]=V(n)$. The self-normalized IS estimator
$\hat{r}_{\mathrm{SNIS}}(n)=\frac{\sum_{\ell}w_{\ell}r_{\ell}}{\sum_{\ell}w_{\ell}}$
is consistent with variance $O(1/\mathrm{ESS})$, where
$\mathrm{ESS}=\frac{(\sum_{\ell}w_{\ell})^{2}}{\sum_{\ell}w_{\ell}^{2}}.$
C.3 Depth-wise Baseline (Control Variates)
For $K$ siblings at depth $i$, let fused returns $\bar{r}^{(b)}$, and group mean $\bar{r}_{i}=\tfrac{1}{K}\sum_{b}\bar{r}^{(b)}$. Define
$A_{i}^{(b)}=\bar{r}^{(b)}-\bar{r}_{i},\qquad\sum_{b}A_{i}^{(b)}=0.$ (10)
Let $g_{i}^{(b)}(\theta)=\nabla_{\theta}\log p_{\theta}(\text{branch }b\text{ at depth }i)$. Then
$\widehat{\nabla J}_{\mathrm{group}}=\frac{1}{K}\sum_{b=1}^{K}A_{i}^{(b)}\,g_{i}^{(b)}$
is an unbiased gradient estimator with strictly smaller variance than $\widehat{\nabla J}_{\mathrm{single}}=\frac{1}{K}\sum_{b}\bar{r}^{(b)}g_{i}^{(b)}$, unless $\mathrm{Cov}(\bar{r}^{(b)},g_{i}^{(b)})=0$.
C.4 Continuous Rewards and Concentration
Assume $r$ is $L$-Lipschitz and the SDE flow $\Psi_{i+1\to N}$ is $K_{i}$-Lipschitz. Then
$|r(z_{N}^{(b)})-r(z_{N}^{(b^{\prime})})|\leq LK_{i}g_{i}\|\xi_{b}-\xi_{b^{\prime}}\|.$
Thus $r(z_{N}^{(b)})$ is sub-Gaussian with parameter $\mathcal{O}(L^{2}K_{i}^{2}g_{i}^{2})$. Averaging over $|L(n)|$ leaves gives
$\Pr\big{(}|\bar{r}(n)-V(n)|\geq\varepsilon\mid n\big{)}\leq 2\exp\left(-c\cdot|L(n)|\,\varepsilon^{2}/(L^{2}K_{i}^{2}g_{i}^{2})\right).$
Appendix D Additional Experiments
D.1 More quantitative result
Table 3 reports more results.
Table 3: Ablation study of BranchGRPO under different design choices. Best and second-best per column are in bold and underline. All results are obtained under the same training setup as in Section 4.3.
[…]
Branch Density
$(0,3,6,9)$ 13.68 13.68 493 0.363 0.229 1.233 0.374
$(0,4,8,12)$ 11.56 11.56 416 0.359 0.229 1.238 0.374
$(0,5,10,15)$ 9.44 9.44 340 0.354 0.228 1.220 0.366
Reward Fusion
Uniform Fusion 13.68 13.68 493 0.361 0.229 1.218 0.368
Path-Weighted Fusion 13.68 13.68 493 0.363 0.229 1.233 0.374
Pruning Strategy
Width Pruning 13.68 8.625 314 0.364 0.230 1.300 0.374
Depth Pruning 13.68 8.625 314 0.369 0.231 1.319 0.381
D.2 More Text2Image Results
Figure 8:
Figure 9:
Figure 10:
Figure 11:
Figure 12:
D.3 More Image2Video Results
(a) Case 1
(b) Case 2
(c) Case 3
(a) Case 4
(b) Case 5
D.4 Failure Cases
Figure 15: Failure case
Figure 16: Failure case